OpenStudy (anonymous):

Solve. 4 t plus 6 is greater than or equal to ten Which graph shows the solution? A number line is numbered zero through ten. A solid circle is over the 1 and an arrow extends from this circle to the right. A number line is numbered zero through ten. A solid circle is over the 4 and an arrow extends from this circle to the right. A number line is numbered ten through twenty. An open circle is over the twelve and an arrow extends from this circle to the right. A number line is numbered ten through twenty. An open circle is over the sixteen and an arrow extends from this circl

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

"4 t plus 6 is greater than or equal to ten" needs to be converted to symbolic form, involving the "equal to or greater than" symbol. Can you do this conversion, Marrero? Which graph shows the solution?

OpenStudy (mathmale):

Try to build on what you learned in the previous problem, OK?

OpenStudy (anonymous):

is it b?

OpenStudy (mathmale):

sorry, no. Would you mind showing all of your work, please? That way i could know exactly how best to help you.

OpenStudy (anonymous):

i just dont understand it

OpenStudy (mathmale):

Translate "4 t plus 6 is greater than or equal to ten" into symbols. What does "4t plus 6" look like in symbols? remember that that "plus" means "addition."

OpenStudy (anonymous):

4t + 6

OpenStudy (mathmale):

Right. And: "4 t plus 6 is greater than or equal to ten." So, 4t + 6 = what?

OpenStudy (anonymous):

the t is still there

OpenStudy (anonymous):

4 + 6 = 10

OpenStudy (mathmale):

Except that you still have that " t " in there. 4 + 6 = 10 NO 4t + 6 = 10 YES subtract 6 from both sides of this equation. Then solve it for t.

OpenStudy (anonymous):

ok it makes 4

OpenStudy (anonymous):

t is 4

OpenStudy (mathmale):

If 4t + 6 = 10, and we subtract 6 from both sides, we get: 4t=4 Divide both sides by 4 (why?)

OpenStudy (anonymous):

idk

OpenStudy (mathmale):

4t = 4 ---- ---- 4 4 Simplify both sides, please.

OpenStudy (anonymous):

it makes 2

OpenStudy (mathmale):

\[\frac{ 4t }{ 4 }=t, so. \frac{ 4 }{ 4 }=?\]

OpenStudy (mathmale):

Remember, treat each side of the equation separately.

OpenStudy (anonymous):

1

OpenStudy (anonymous):

4/4 = 1

OpenStudy (mathmale):

Right, and so (4t/4) = t = 1. Now, because the original problem statement said "greater than or equal to," write\[t \ge 1\] and also determine which of the four possible asnwers is correct; it is not B.

OpenStudy (anonymous):

ts a thx

OpenStudy (mathmale):

Good for you! While i will not be on the Internet much longer, I'll look forward to helping you again in the future. Best to you, Javier.